:An attractive force of 12kN is experienced by two charges of magnitude 4nC and 7nC. At what

distance must these two charges be to experience this force?


1
Expert's answer
2022-03-07T12:21:23-0500

Answer

Distance

"r=\sqrt{\frac{9*10^9*4*7*10^{-18}}{12000}}\\=4.5*10^{-6}m"